OpenStudy (anonymous):

One pipe can fill a tank in 36 minutes, a second can fill it in 9 minutes, and a third can fill it in 12 minutes. If the tank is empty, how long will the three pipes, operating together, take to fill it?

OpenStudy (anonymous):

Solve the following for x:\[\frac{1}{36}+\frac{1}{9}+\frac{1}{12}=\frac{1}{x} \]x = 4.5 minutes